OpenStudy (anonymous):

Find the slope of the line passing through (5, -6) and (4, 3).

OpenStudy (anonymous):

3-(-6)/4-5=-9

OpenStudy (anonymous):

-1/9 fraction negative gradient

OpenStudy (anonymous):

\[\frac{-6-3}{5-4}=-9\]